OpenStudy (anonymous):

Write the equation of a line that passes through the points (-8, 8) and (0, 1).

OpenStudy (anonymous):

To solve this problem .... use the following steps: Get the slope of the line: \[\frac{ Y2 - Y1 }{ X2 - X1 }\] Your points are: (-8,8) and (0,1) \[\frac{ 1 - 8 }{ 0 - (-8) } = \frac{ -7 }{ 8 }\] Thats your slope. Take your slope and plug it into the following formula: Y=MX+B \[Y = MX + B ....... y = \frac{ -7 }{ 8 }x + b\] Now your going to take one of the points .... and plug them in to get the value of b. For this one ... ill use the easy point (0,1) X = 0 .... and Y = 1 Plug your points into the formula Y = MX + B Remember (MX) is your slope! Your problem should look like this: \[1 = \frac{ -7 }{ 8 }(0) + b\] \[1 = 0 + b\] \[1 = b\] Now you have all your numbers.Plug them up and you have your equation of a line that passes through the two points: \[y=\frac{ -7 }{ 8 }x+1\]